OpenStudy (anonymous):

find the product (2x+5)^2

OpenStudy (anonymous):

4x^2+20x+25 btw wolframalpha.com can also do this

OpenStudy (anonymous):

(2x+5)(2x+5)

OpenStudy (anonymous):

Multiply through.

OpenStudy (anonymous):

i see.....5^2=25

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

5*5 = (5)^2

OpenStudy (anonymous):

2x^2=2x.2x 5^2=5.5

OpenStudy (anonymous):

so the answer is (2x+5)(2x+5) @olympics

OpenStudy (anonymous):

Nope, you have to multiply it now.

OpenStudy (anonymous):

(2x+5)(2x+5) = 4x^2 +10x+10x+25 collect like terms, factor, etc.

OpenStudy (anonymous):

4x^2+10x+10x+25 4^2+20x+25

OpenStudy (anonymous):

Now factor!

OpenStudy (anonymous):

can't think of two numbers which product is 25 and sum 20

OpenStudy (anonymous):

x = -5/2 would be your solution

OpenStudy (anonymous):

If you set to 0 you don't have to do any of that stuff...

OpenStudy (anonymous):

how @iambatman

OpenStudy (anonymous):

use quadratic formula u'll get x= -5/2

OpenStudy (anonymous):

Olympics gave expanded form, you just need the root I assume.

OpenStudy (anonymous):